WHat is the length and width of a rectangular box of it is volume is 240cm3 and its heigth is 15
Found 2 solutions by CubeyThePenguin, greenestamps:
Answer by CubeyThePenguin(3113)

length = L
width = W
height = H = 15

Volume = LWH = 240

Area of base = WH = 240/15 = 16

Answer by greenestamps(13367)



No solution -- there is not enough information.

The volume is the area of the base, multiplied by the height.

Given that the volume is 240 and the height is 15, we can conclude that the area of the base is 240/15 = 16.

But there is not enough information to be able to determine the length and width.